OpenStudy (anonymous):

f(x)= -√x=1 Graph the functions and describe the domain and range of each function. anybody?

hero (hero):

your function has two equal signs in it.

OpenStudy (anonymous):

f(x)= -√x+1

hero (hero):

domain x >=-1 range y <= 0

OpenStudy (anonymous):

If you take the square root of a negative number, it is not a member of the real number system. x cannot be negative. The domain of f(x) is \[[0,\infty)\]

hero (hero):

^not true, x can be negative

OpenStudy (anonymous):

Yes, but it will not yield a real number

OpenStudy (anonymous):

I'm assuming he wants real numbers

hero (hero):

x + 1 cannot be less than zero

OpenStudy (anonymous):

Umm, how about I just post the thing that has all these on them

OpenStudy (anonymous):

It says \[-\sqrt{x}=1\]

OpenStudy (anonymous):

How can x be negative?

hero (hero):

sqrt(x+1) not sqrt(x)

OpenStudy (anonymous):

Oh i was going by the original post. Didn't see the updated one.

OpenStudy (anonymous):

\[f(x)=-\sqrt{x+1} \,\,\text {or} f(x)=-\sqrt{x}+1?\]

OpenStudy (anonymous):

That's what I thought cinar

OpenStudy (anonymous):

that was a typo. sorry you guys!

OpenStudy (anonymous):

So, which is it?

hero (hero):

Use parentheses

OpenStudy (anonymous):

f(x)= -√x+1

OpenStudy (anonymous):

so 1 is not under the square root?

OpenStudy (anonymous):

it is. x+1 are under the square root

OpenStudy (anonymous):

OK, use parentheses next time \[\sqrt{x+1}>0, x>=-1\] Domain is \[[-1,\infty)\]

OpenStudy (anonymous):

then domain is [-1,infinite) range is [0, infinite)

OpenStudy (anonymous):

thank you!